Two boats leave port at the same time, one boat sails due north and the other due west. when they are 80 miles apart, the boat traveling north has sailed 16 miles farther than the boat sailing west. Find the distance each boat has traveled.
Answer by richwmiller(17219)

Two boats leave port at the same time, one boat sails due north and the other due west. when they are 80 miles apart, the boat traveling north has sailed 16 miles farther than the boat sailing west. Find the distance each boat has traveled.
x^2+(x+16)^2=80^2
x=48 miles
x+16=64 miles